A person takes a loan of Rs.5000 at compound interest compounded annually. He pays Rs.2500 at the end of first year and Rs.3472 at the end of second year and clears the debt. If the rate of interest is the same in both the years. Find the rate of interest?
Answer by addingup(3677)

(5000*(1+r)-2500)+r(5000*(1+r)-2500)= 3472 Let me rewrite it like this:
-2500+5000(r+1)+r(5000(r+1)-2500)= 3472 Now we'll write the left side as a quadratic:
5000r^2+7500r+2500= 3472 divide both sides by 5000
r^2+(3r/2)+1/2= 434/625 Subtract 1/2 from both sides:
r^2+(3r/2)= 243/1250 Add 1/2 of the coefficient of r squared to both sides
r^2+(3r/2)+9/16= 7569/10000 Factor the left
(r+3/4)^2= 7569/10000 Take the square of both sides to get rid of the exponent
r+3/4= 87/100 or r+3/4= -87/100 Subtract 3/4 on all four sides:
r= 3/25 or r= -81/50
Since we are looking for a positive number, let's take 3/25:
3/25= 0.12 (12%) is your answer














0= 3472
(5000*1.12)-2500+.12(5000*1.12)